(1 point) Determine the number of regular semi annual payments for a $17000 loan if the interest rate is J2-11 .8% and the payments are: a)Semi annual payments of $3600 b)Semi annual payments of $4300: c)Semi annual payments of $2200:
Answer #1

1.

=NPER(11.8%/2,-3600,17000)=5.70 or 5 regular

2.

=NPER(11.8%/2,-4300,17000)=4.63 or 4 regular

3.

=NPER(11.8%/2,-2200,17000)=10.62 or 10 regular